OpenStudy (rainbow_dash):

Identify a1 in a geometric series for which S8 = 42.5 and r = -0.5.

OpenStudy (anonymous):

\(s_8=s_1\times r^7\)

OpenStudy (rainbow_dash):

ohhh raised to the 7. i didn't know what to do because of the lack of an n :3

OpenStudy (anonymous):

so \[s_1=\frac{s_8}{r^7}\] in your case \(s_1=\frac{42.5}{(-.5)^7}\)

OpenStudy (anonymous):

\(s_1,s_2,s_3,s_4,...\) \(s_1,s_1r,s_2r^2,s_3r^3,...\)

OpenStudy (rainbow_dash):

that can't be right...i got -6071 for the -.5^7, but the highest possible answer is 63.75

OpenStudy (anonymous):

you put it in the calculator wrong \((-.5)^7\) is a little tiny number

OpenStudy (rainbow_dash):

i meant 42.5/that tiny number :( my mistake

OpenStudy (anonymous):

maybe i read the problem wrong maybe it is \(\sum_{k=1}^8a_k=42.5\)

OpenStudy (rainbow_dash):

oh god that sigma is scary.

OpenStudy (anonymous):

it is the 8th partial sum maybe

OpenStudy (rainbow_dash):

sorry but what does this mean?

OpenStudy (anonymous):

\[a_1+a_2+a_3+a_4+a_5+a_6+a_7+a_8=42.5\]

OpenStudy (anonymous):

formula is \[\frac{1-a^{n+1}}{1-r}\] so you get \[\frac{1-a_1(-.5)^8}{1+.5}=42.5\]

OpenStudy (anonymous):

solve this for \(a_1\)

OpenStudy (anonymous):

oh damn that is wrong, first term is not 1, sorry \[\frac{a_1(1-a_1(-.5)^8)}{1-(-.5)}=42.5\]

OpenStudy (anonymous):

\[S_n = \frac{a(1-r^n)}{1-r}\]

OpenStudy (anonymous):

That a is \(a_1\)